OpenStudy (anonymous):

Given the binomials (x - 1), (x - 3), (x + 3), and (x + 5), which one is a factor of f(x) = x^3 + 6x^2 + 12x + 35?

OpenStudy (anonymous):

(x - 1) (x - 3) (x + 3) (x + 5)

ganeshie8 (ganeshie8):

use factor theorem + synthetic division

ganeshie8 (ganeshie8):

if x - 1 is a factor, then the remainder in division will be 0

OpenStudy (anonymous):

\[f(x) = x^3 + 6x^2 + 12x + 35? \] to make it easier to read

ganeshie8 (ganeshie8):

u familiar wid synthetic diviosn right ?

OpenStudy (anonymous):

somewhat

ganeshie8 (ganeshie8):

setup the division

ganeshie8 (ganeshie8):

lets test first if x-1 is a factor 1 | 1 6 12 35 | --------------------------

ganeshie8 (ganeshie8):

do u knw the steps ?

OpenStudy (anonymous):

not much... I'm guessing divide the terms by -1?

ganeshie8 (ganeshie8):

nope, just move the first coefficient 1 down. make sure u understand the steps cuz u gona test remaining factors :)

ganeshie8 (ganeshie8):

1 | 1 6 12 35 | -------------------------- 1

ganeshie8 (ganeshie8):

next, multiply this 1 wid the 1 on left and write it down the next coefficient

ganeshie8 (ganeshie8):

1 | 1 6 12 35 | 1 -------------------------- 1

ganeshie8 (ganeshie8):

next add 6 and 1 and write it below

ganeshie8 (ganeshie8):

1 | 1 6 12 35 | 1 -------------------------- 1 7

ganeshie8 (ganeshie8):

repeat the steps :- miltiply 7 and 1, write it below next coeffecient

ganeshie8 (ganeshie8):

1 | 1 6 12 35 | 1 7 -------------------------- 1 7

ganeshie8 (ganeshie8):

add, multiply, add multiply..... keep going

OpenStudy (anonymous):

uhh 1* 12 = 12... 12 + 7 = 19?

ganeshie8 (ganeshie8):

1 | 1 6 12 35 | 1 7 19 -------------------------- 1 7 19 |54 so remainder is 54, which is not 0. so x-1 is not a factor.

ganeshie8 (ganeshie8):

similarly, test next factor

ganeshie8 (ganeshie8):

u need to have lots of patience ok -_-

ganeshie8 (ganeshie8):

lets test first if x-3 is a factor 3 | 1 6 12 35 | --------------------------

OpenStudy (anonymous):

I see so a lot of testing lol

ganeshie8 (ganeshie8):

give it a try

ganeshie8 (ganeshie8):

yes after this, two more.

OpenStudy (anonymous):

so we want 0 as our remainder?

ganeshie8 (ganeshie8):

yes, if u get 0 as remainder, then its a factor.

OpenStudy (anonymous):

3 * 1 = 3, 3 + 1 = 4 3 * 6 = 18, 18 + 4 = 22 or is it 6 + 4 = 10?

ganeshie8 (ganeshie8):

i want u do the synthetic division *properly* in the synthetic division looking format :)

OpenStudy (anonymous):

lol

ganeshie8 (ganeshie8):

lets test first if x-3 is a factor 3 | 1 6 12 35 | 3 27 117 ------------------------- 1 9 39 | 152 here remainder is 152, which is not 0. so x-3 is not a factor.

ganeshie8 (ganeshie8):

test next guy

ganeshie8 (ganeshie8):

lets test first if x+3 is a factor -3 | 1 6 12 35 | -------------------------- 1

ganeshie8 (ganeshie8):

see if u can do this

ganeshie8 (ganeshie8):

(in synthetic division format :) )

ganeshie8 (ganeshie8):

u gave up ?

OpenStudy (anonymous):

-3| 1 6 12 35 | -3 -18-216 --------------------------- -3 -18-216-7560 lol?

OpenStudy (anonymous):

I feel like I did something wrong when I ended up with -7560

ganeshie8 (ganeshie8):

NO ! u need to add. not multiply

OpenStudy (anonymous):

Right...

ganeshie8 (ganeshie8):

-3| 1 6 12 35 | -3 --------------------------- -3 3

ganeshie8 (ganeshie8):

do it again

ganeshie8 (ganeshie8):

now multiply 3 wid left side -3

ganeshie8 (ganeshie8):

and write it below 12

OpenStudy (anonymous):

-3|1 6 12 35 | -3 3 15 ---------------------- -3 3 15 40?

ganeshie8 (ganeshie8):

sir, 3*-3 = -9

OpenStudy (anonymous):

ah I'm so confused lol

ganeshie8 (ganeshie8):

-3|1 6 12 35 | -3 -9 ---------------------- 1 3

ganeshie8 (ganeshie8):

add 12 and -9 write the result below

OpenStudy (anonymous):

3

ganeshie8 (ganeshie8):

-3|1 6 12 35 | -3 -9 ---------------------- 1 3 3

ganeshie8 (ganeshie8):

multiply this 3 wid left side -3 and write the result below next coefficient.

ganeshie8 (ganeshie8):

-3|1 6 12 35 | -3 -9 -9 ---------------------- 1 3 3

ganeshie8 (ganeshie8):

add both

ganeshie8 (ganeshie8):

-3|1 6 12 35 | -3 -9 -9 ---------------------- 1 3 3 | 26 the remainder is 26, which is not 0 so, x+3 is not a factor.

ganeshie8 (ganeshie8):

test the last one.

OpenStudy (anonymous):

+ 5 is our last one right?

ganeshie8 (ganeshie8):

x+5 so u need to put -5 in synthetic division

ganeshie8 (ganeshie8):

-5|1 6 12 35 | ---------------------- 1

ganeshie8 (ganeshie8):

carry out same steps, u should get 0 this time. give ur best try :)

OpenStudy (anonymous):

alright so first I would multiply 1 * -5 right?

ganeshie8 (ganeshie8):

yes

OpenStudy (anonymous):

so -5 + 6 now?

OpenStudy (anonymous):

which is 1

ganeshie8 (ganeshie8):

yes, synthetic division format please -.-

OpenStudy (anonymous):

man that takes forever to draw out lol

ganeshie8 (ganeshie8):

copy paste ?

ganeshie8 (ganeshie8):

-5|1 6 12 35 | ---------------------- 1

ganeshie8 (ganeshie8):

-5|1 6 12 35 | ---------------------- 1

ganeshie8 (ganeshie8):

see it wont take any time :) copy paste the thing and fill the second and 3rd rows ok

ganeshie8 (ganeshie8):

-5|1 6 12 35 | -5 -5 -35 ---------------------- 1 1 7 | 0 since 0 is the remainder here, x+5 is a factor.

OpenStudy (anonymous):

sorry about that had to answer the phone

OpenStudy (anonymous):

ohh ok I see how you did that

OpenStudy (anonymous):

Thanks for the help man! Can finally go to the gym now :D

ganeshie8 (ganeshie8):

Enjoy the workout cya !